OpenStudy (anonymous):

Given the function f(x) = 5x, Section A is from x = 0 to x = 1 and Section B is from x = 2 to x = 3. Part A: Find the average rate of change of each section. (4 points) Part B: How many times greater is the average rate of change of Section B than Section A? Explain why one rate of change is greater than the other. (6 points)

OpenStudy (anonymous):

please help i will fan

OpenStudy (ryuga):

medal + testmonial

OpenStudy (anonymous):

i don't have any metals or i would give you one

OpenStudy (anonymous):

i will give you a testimonial, what do you want it to say

OpenStudy (ryuga):

The question is defective, or at least is trying to lead you down the primrose path. The function is linear, so the rate of change is the same no matter what interval (section) of it you're looking at. The "rate of change" is just the slope of the function in the section. That's (change in f(x) ) / (change in 'x') between the ends of the section. In Section A: Length of the section = (1 - 0) = 1 f(1) = 5 f(0) = 0 change in the value of the function = (5 - 0) = 5 Rate of change = (change in the value of the function) / (size of the section) = 5/1 = 5 In Section B: Length of the section = (3 - 2) = 1 f(3) = 15 f(2) = 10 change in the value of the function = (15 - 10) = 5 Rate of change = (change in the value of the function) / (size of the section) = 5/1 = 5 Part A: The average rate of change of each section is 5. Part B: The average rate of change of Section B is equal to the average rate of change of Section A. Explanation: The average rates of change in every section are equal because the function is linear, its graph is a straight line, and the rate of change is just the slope of the graph.

OpenStudy (ryuga):

medal = best response

OpenStudy (ryuga):

@mathmale did i do a good job of explaining?

OpenStudy (anonymous):

Yess so this is my answer? right In Section A: Length of the section = (1 - 0) = 1 f(1) = 5 f(0) = 0 change in the value of the function = (5 - 0) = 5 Rate of change = (change in the value of the function) / (size of the section) = 5/1 = 5 In Section B: Length of the section = (3 - 2) = 1 f(3) = 15 f(2) = 10 change in the value of the function = (15 - 10) = 5 Rate of change = (change in the value of the function) / (size of the section) = 5/1 = 5 Part A: The average rate of change of each section is 5. Part B: The average rate of change of Section B is equal to the average rate of change of Section A. Explanation: The average rates of change in every section are equal because the function is linear, its graph is a straight line, and the rate of change is just the slope of the graph.

OpenStudy (anonymous):

I have one more question!!? @Ryuga

OpenStudy (ryuga):

you are corerect

OpenStudy (anonymous):

can you help me with one more question???? @Ryuga

OpenStudy (ryuga):

close this and ask again

OpenStudy (anonymous):

kk